OpenStudy (anonymous):

(-5 + x)(-3 + y)(3-3)

OpenStudy (anonymous):

3-3=0 so (-5+x)(-3+y)(3-3)=0

OpenStudy (anonymous):

(-5+x)(-3+y)(3-3)=0

OpenStudy (anonymous):

(3-3)(-5+x)(-3+y)=0 why? 3-3=0

OpenStudy (anonymous):

three out of three equals zero!